### v1

# On the local thermodynamic relations in relativistic spin hydrodynamics

Abstract
We demonstrate, by providing two specific examples, that the local differential thermodynamic relations used as educated guesses in relativistic hydrodynamics with spin, do not hold even at global thermodynamic equilibrium, where the expression of the entropy current is totally unambiguous. We show, by using a rigorous quantum statistical method, that for massless free fermions and massive free fermions with rotation and acceleration at global thermodynamic equilibrium, the derivative of the pressure function with respect to the spin potential differs from the spin density and acquires a correction of the same order. Therefore, for an accurate determination of the constitutive relations in relativistic spin hydrodynamics, the traditional method of assuming differential thermodynamic relations is not appropriate.
preprint:
I Introduction
Prompted by the measurements of spin polarization of particles produced in relativistic heavy-ion collisions [1, 2, 3, 4], there has been a surge in the development of relativistic hydrodynamic frameworks that incorporate spin as an additional hydrodynamic degree of freedom [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22].
Relativistic spin hydrodynamics, in essence, includes the conservation of the angular momentum-boost tensor besides that of energy-momentum. The operator angular momentum-boost density reads:
$\widehat{\cal J}^{\lambda\mu\nu}=x^{\mu}{\widehat{T}}^{\lambda\nu}-x^{\nu}{\widehat{T}}^{\lambda\mu}+{\widehat{\cal S}}^{\lambda\mu\nu}\equiv\widehat{\cal L}^{\lambda\mu\nu}+{\widehat{\cal S}}^{\lambda\mu\nu}\;,$ (1)
where ^$\widehat{\cal L}^{\lambda\mu\nu}$ is the so-called orbital part and ${\widehat{\cal S}}^{\lambda\mu\nu}$ is the spin tensor. Due to the conservation of $\widehat{\cal J}^{\lambda\mu\nu}$, the divergence of the spin tensor equals the anti-symmetric part of the energy-momentum tensor
$\partial_{\lambda}{\widehat{\cal S}}^{\lambda\mu\nu}={\widehat{T}}^{\nu\mu}-{\widehat{T}}^{\mu\nu}\,.$ (2)
As it is well known, the stress-energy and spin tensors are not uniquely defined and one may obtain different pairs of those tensors through a so-called pseudo-gauge transformation [23, 24, 25]
$\hat{T}^{\prime}{{}^{\mu\nu}}$ $=$ $\hat{T}^{\mu\nu}+\frac{1}{2}\partial_{\lambda}(\hat{\Pi}^{\lambda,\mu\nu}+\hat{\Pi}^{\nu,\mu\lambda}+\hat{\Pi}^{\mu,\nu\lambda})\,,$ (3)
$\hat{\mathcal{S}}^{\prime}{{}^{\lambda,\mu\nu}}$ $=$ $\hat{\mathcal{S}}^{\lambda,\mu\nu}-\hat{\Pi}^{\lambda,\mu\nu}+\partial_{\rho}\hat{\Theta}^{\mu\nu,\lambda\rho}\,,$ (4)
where $\hat{\Pi}$ and $\hat{\Theta}$ are known as super-potentials satisfying, $\hat{\Pi}^{\lambda,\mu\nu}=-\hat{\Pi}^{\lambda,\nu\mu}$ and $\hat{\Theta}^{\mu\nu,\lambda\rho}=-\hat{\Theta}^{\nu\mu,\lambda\rho}=-\hat{\Theta}^{\mu\nu,\rho\lambda}$, respectively. The new pair of stress-energy and spin tensors, however, satisfy the same dynamical equations and do not change the integrated conserved charges. However, both the entropy and entropy production rate are apparently dependent on the choice of the pseudo-gauge [26], thus making the study of the entropy current in spin hydrodynamics an important subject.
Requiring the local entropy production rate to be positive definite, there have been numerous derivations of the constitutive relations of the spin tensor and the stress-energy tensor as a function of the gradients of the thermo-hydrodynamic fields [9, 12, 27, 13, 28, 16, 22, 29, 30]. In those formulations, the entropy current is obtained by assuming the validity of the thermodynamic relations:
$Ts$ $=$ $\rho+p-\mu n-\frac{1}{2}\omega_{\mu\nu}S^{\mu\nu}\,\,,$ (5)
${\rm d}p$ $=$ $s\,{\rm d}T+n\,{\rm d}\mu+\frac{1}{2}S^{\mu\nu}{\rm d}\omega_{\mu\nu}\,,$ (6)
with $s$, $n$, $\rho$, $p$, and $S^{\mu\nu}\equiv{\cal S}^{\lambda,\mu\nu}u_{\lambda}$ being the entropy density, charge density, energy density, pressure, and spin density, respectively, whereas $T$, $\mu$, and $\omega_{\mu\nu}$ are temperature, chemical potential, and spin potential, respectively  [^1] In this work, we will name the quantity $\omega=\Omega\,T$ as spin potential, whereas $\Omega$ will be referred to as reduced spin potential. Note that in previous literature, $\Omega$ is referred to as either spin potential or spin chemical potential, but these terms do not accord well to the usual naming of $\mu$ as chemical potential. The ratio $\mu/T=\zeta$ will be referred to as reduced chemical potential..
However, it was argued in ref. [26] that the differential relation (6) does not generally apply, so that it is not suitable to derive the entropy production rate, whence the constitutive equations of relativistic spin hydrodynamics at the needed level of precision. It was shown that while it is possible to define pressure such that the (5) holds, as well as $\partial p/\partial T|_{\mu,\omega}=s$, it is not guaranteed that:
$\frac{\partial p}{\partial\mu}\Big{|}_{T,\omega}$ $=$ $n\,,$ (7)
$\frac{\partial p}{\partial\omega_{\lambda\nu}}\Big{|}_{T,\mu}$ $=$ $S^{\lambda\nu}\,.$ (8)
[…]
$T\frac{\partial s}{\partial\omega_{\mu\nu}}\Big{|}_{T,\mu}-\frac{\partial\rho}{\partial\omega_{\mu\nu}}\Big{|}_{T,\mu}+\mu\frac{\partial n}{\partial\omega_{\mu\nu}}\Big{|}_{T,\mu}+\frac{1}{2}\omega_{\rho\sigma}\frac{\partial S^{\rho\sigma}}{\partial\omega_{\mu\nu}}\Big{|}_{T,\mu}=0,$
which is not guaranteed. The same conclusion applies if $s$ was defined through the (5) and $p$ was known. The only way to check the validity of the above equation, or tantamount (8), is to calculate the thermodynamic functions at the second order in vorticity within a quantum statistical framework.
In this paper, we will do this and we will provide two specific examples of the breaking of the relation (8): for a system of massless free fermions at global thermodynamic equilibrium with rotation and acceleration, for which we know the exact solution; and for a system of massive free fermions at global thermodynamic equilibrium with rotation and acceleration, for which we know the expression of the thermodynamic functions up to second order in vorticity. Since these results are obtained at global thermodynamic equilibrium, where all thermodynamic quantities are unambiguous and no freedom in defining entropy current can be invoked, we conclude that the relation (6) is incorrect, with all due implications.
II Local equilibrium density operator and the entropy current
To make this article self-contained, we will first briefly review the derivation of the entropy current presented in [26].
For a relativistic fluid, by applying the quantum-statistical method, one obtains the local equilibrium density operator, ${\widehat{\rho}}_{LE}$, by maximizing the entropy:
$S=-{\rm Tr}\left({\widehat{\rho}}\log{\widehat{\rho}}\right)\,,$ (9)
over some space-like hypersurface $\Sigma$ with normal unit vector $n$ constraining the average values of the energy density, momentum density, charge density, and spin density to be the same as their original values:
$n_{\mu}T^{\mu\nu}=n_{\mu}{\rm Tr}\left({\widehat{\rho}}\,{\widehat{T}}^{\mu\nu}\right)\,,$ $n_{\mu}j^{\mu}=n_{\mu}{\rm Tr}\left({\widehat{\rho}}\,{\widehat{j}}^{\mu}\right)\,,$
$n_{\mu}{\cal S}^{\mu\lambda\nu}$ $=$ $n_{\mu}{\rm Tr}\left({\widehat{\rho}}\,{\widehat{\cal S}}^{\mu\lambda\nu}\right)\,,$ (10)
where ${\widehat{T}}^{\mu\nu}$, ${\widehat{\cal S}}^{\mu\lambda\nu}$ are the energy-momentum and spin tensor operators and ${\widehat{j}}^{\mu}$ is the vector-current operator. The density operator which solves the problem is the local equilibrium density operator:
${\widehat{\rho}}_{\rm LE}=\frac{1}{Z_{\rm LE}}\exp\left[-\hat{\Upsilon}\right]\,,$ (11)
with:
$\hat{\Upsilon}=\int_{\Sigma}{\rm d}\Sigma_{\mu}\,\left({\widehat{T}}^{\mu\nu}\beta_{\nu}-\zeta{\widehat{j}}^{\mu}-\frac{1}{2}\Omega_{\lambda\nu}{\widehat{\cal S}}^{\mu\lambda\nu}\right)\,.$ (12)
In the equation (12) $\beta_{\nu}$, $\zeta$ and $\Omega_{\lambda\nu}$ are the Lagrange multiplier fields defined, respectively, as the four-temperature vector, the reduced chemical potential and the reduced spin potential:
$\beta^{\mu}=\frac{u^{\mu}}{T}\,,\qquad\zeta=\frac{\mu}{T}\,,\qquad\Omega^{\lambda\nu}=\frac{\omega^{\lambda\nu}}{T}\,,$ (13)
so that $T=1/\sqrt{\beta^{2}}$ is the proper temperature. At the global thermodynamic equilibrium, $\beta_{\mu}$ is a Killing vector, $\Omega^{\mu\nu}$ is constant and equals thermal vorticity $\varpi^{\mu\nu}$ and $\zeta$ is constant:
$\beta_{\mu}$ $=$ $b_{\mu}+\varpi_{\mu\nu}x^{\nu}\,,\quad{\rm with}\quad b_{\mu}={\rm const.}\,,$ (14)
$\Omega_{\mu\nu}$ $=$ $\varpi_{\mu\nu}=-\frac{1}{2}\left(\partial_{\mu}\beta_{\nu}-\partial_{\nu}\beta_{\mu}\right)={\rm const.}\,,$ (15)
$\zeta$ $=$ ${\rm const.}\,,$ (16)
and in this case the local equilibrium density operator becomes:
${\widehat{\rho}}_{\rm GE}=\frac{1}{Z_{\rm GE}}\exp\left[-b\cdot{\widehat{P}}+\frac{1}{2}\varpi:{\widehat{J}}+\zeta{\widehat{Q}}\right]\,,$ (17)
where the subscript ‘GE’ stands for global equilibrium. There are other useful relations at global equilibrium involving the acceleration field $a^{\mu}$ and vorticity field $\omega^{\mu}$:
$\alpha^{\mu}$$\equiv$ $\varpi^{\mu\nu}u_{\nu}=\frac{1}{T}u^{\nu}\partial_{\nu}u^{\mu}=\frac{a^{\mu}}{T}\,,$ (18)
$w^{\mu}$$\equiv$ $-\frac{1}{2}\epsilon^{\mu\nu\rho\sigma}\varpi_{\nu\rho}u_{\sigma}=-\frac{1}{2T}\epsilon^{\mu\nu\rho\sigma}\partial_{\rho}u_{\nu}u_{\sigma}=\frac{\omega^{\mu}}{T}\,,$
hence, since at global equilibrium $\varpi_{\mu\nu}=\Omega_{\mu\nu}=\omega_{\mu\nu}/T$, we have that the spin potential equals the kinematic vorticity:
$\frac{a^{\mu}}{T}=\frac{1}{T}\omega^{\mu\nu}u_{\nu}\,,\qquad\qquad\frac{\omega^{\mu}}{T}=-\frac{1}{2T}\epsilon^{\mu\nu\rho\sigma}\omega_{\nu\rho}u_{\sigma}\,,$
$\implies\omega_{\mu\nu}=a_{\mu}u_{\nu}-a_{\nu}u_{\mu}+\epsilon_{\mu\nu\rho\sigma}\omega^{\rho}u^{\sigma}\,.$ (19)
It is convenient to define a normalized mean value of operators at local equilibrium:
$\mathcal{X}_{\rm LE}\equiv{\rm Tr}\left({\widehat{\rho}}_{\rm LE}\widehat{\mathcal{X}}\right)-\langle 0|\widehat{\mathcal{X}}|0\rangle\,,$ (20)
where the state $|0\rangle$ in Eq. (20) is the (supposedly non-degenerate) lowest lying state of the operator $\widehat{\Upsilon}$ in the equation (12). Note that this state does not necessarily coincide with the Minkowski vacuum $|0_{M}\rangle$; for instance, if:
$\widehat{\Upsilon}=\frac{1}{T_{0}}\left(\widehat{H}-\mu\widehat{Q}\right)\,,$
for the free Dirac field, the state $|0\rangle$ is the Fermi gas at zero temperature. Indeed, if we take the limit $T\to 0$ in the equation (11) we should recover the state $|0\rangle$:
$\lim_{T\to 0}{\widehat{\rho}_{\rm LE}}=|0\rangle\langle 0|\,,$ (21)
implying that the local equilibrium values defined above must fulfill the following limit:
$\lim_{T\to 0}\mathcal{X}_{\rm LE}=0\,.$ (22)
The first step in order to derive the entropy current is to prove that the logarithm of the partition function is extensive, meaning it can be written as an integral of some four-vector field $\phi^{\mu}$ over the local equilibrium hypersurface $\Sigma$ as:
$\log Z_{\rm LE}=\int_{\Sigma}{\rm d}\Sigma_{\mu}\;\phi^{\mu}-\langle 0|\widehat{\Upsilon}|0\rangle\,.$ (23)
The field $\phi^{\mu}$ is defined as the thermodynamic potential current [33] and can be calculated as follows [31, 26]:
$\phi^{\mu}(x)=\int_{0}^{T(x)}\frac{{\rm d}T^{\prime}}{T^{\prime 2}}\;\Big{(}u_{\nu}(x)T^{\mu\nu}_{\rm LE}(x)[T^{\prime},\mu,\omega]$ (24)
$-\mu(x)j^{\mu}_{\rm LE}(x)[T^{\prime},\mu,\omega]-\frac{1}{2}\omega_{\lambda\nu}(x){\cal S}^{\mu\lambda\nu}_{\rm LE}(x)[T^{\prime},\mu,\omega]\Big{)}\,.$
In the above equation, the square brackets denote a functional dependence on the thermodynamic fields $T^{\prime},\mu,\omega$. We are now in a position to obtain the entropy current. Starting with Eq. (11) and using Eqs. (20) and (23) we have:
$S$ $=$ $-{\rm Tr}({\widehat{\rho}}_{\rm LE}\log{\widehat{\rho}}_{\rm LE})\,,$ (25)
$=$ $\log Z_{\rm LE}+\int_{\Sigma}{\rm d}\Sigma_{\mu}\;\Big{(}{\rm Tr}({\widehat{\rho}_{\rm LE}}{\widehat{T}}^{\mu\nu})\beta_{\nu}-\zeta{\rm Tr}({\widehat{\rho}_{\rm LE}}{\widehat{j}}^{\mu})$
$-\frac{1}{2}\Omega_{\lambda\nu}{\rm Tr}({\widehat{\rho}_{\rm LE}}{\widehat{\cal S}}^{\mu\lambda\nu})\Big{)}\,,$
$=$ $\int_{\Sigma}{\rm d}\Sigma_{\mu}\;\left(\phi^{\mu}+T^{\mu\nu}_{\rm LE}\beta_{\nu}-\zeta j^{\mu}_{\rm LE}-\frac{1}{2}\Omega_{\lambda\nu}{\cal S}^{\mu\lambda\nu}_{\rm LE}\right)$
$=$ $\int_{\Sigma}{\rm d}\Sigma_{\mu}\;s^{\mu}\,,$
whence we can define the entropy current as:
$s^{\mu}=\phi^{\mu}+T^{\mu\nu}_{\rm LE}\beta_{\nu}-\zeta j^{\mu}_{\rm LE}-\frac{1}{2}\Omega_{\lambda\nu}{\cal S}^{\mu\lambda\nu}_{\rm LE}\,.$ (26)
Because of the constraints (II), the entropy current can be defined as well as:
$s^{\mu}=\phi^{\mu}+T^{\mu\nu}\beta_{\nu}-\zeta j^{\mu}-\frac{1}{2}\Omega_{\lambda\nu}{\cal S}^{\mu\lambda\nu}\,,$ (27)
where $\phi^{\mu}$ is then written as:
$\phi^{\mu}=\int_{0}^{T}\frac{{\rm d}T^{\prime}}{T^{\prime 2}}\;\left(T^{\mu\nu}[T^{\prime}]u_{\nu}-\mu j^{\mu}[T^{\prime}]-\frac{1}{2}\omega_{\lambda\nu}{\cal S}^{\mu\lambda\nu}[T^{\prime}]\right)\,.$ (28)
The form (27) of the entropy current is to be preferred in an out-of-equilibrium situation with respect to (26) [26]. Nevertheless, at global thermodynamic equilibrium the two definitions coincide and there is no ambiguity in its identification. It is worth pointing out that the entropy current, by virtue of the definition (26), vanishes in the limit $T\to 0$:
$\lim_{T\to 0}s^{\mu}=0\,,$
as it should because the entropy is zero when ${\widehat{\rho}_{\rm LE}}$ becomes the pure state $|0\rangle$, see equation (21). This condition entails that, as also implied by equation (II), the mean values of the conserved currents $T^{\mu\nu}$, $j^{\mu}$ etc. vanish in the same limit. In other words, it is understood that in the formulae (27), (28) they are normalized so as to make their value vanishing at zero temperature. Therefore, they do not necessarily vanish in the Minkowski vacuum, as implied by the above discussion.
III Local thermodynamic relations for a free Dirac field at global equilibrium
To obtain the local thermodynamic relations, one requires to contract the entropy current (27) with the fluid four-velocity, $u^{\mu}=\beta^{\mu}/\sqrt{\beta^{2}}=T\beta^{\mu}$ (normalized as $u\cdot u=1$)  [^2] Again, we stress that no ambiguity exists in the definition of a four-velocity at global thermodynamic equilibrium, where $\beta$ is a Killing vector.
$s\equiv s^{\mu}u_{\mu}$ $=$ $\phi^{\mu}u_{\mu}+T^{\mu\nu}\beta_{\nu}u_{\mu}-\zeta j^{\mu}u_{\mu}-\frac{1}{2}u_{\mu}\Omega_{\lambda\nu}{\cal S}^{\mu\lambda\nu}$ (29)
$=$ $\phi\cdot u+\frac{1}{T}\rho-\frac{\mu}{T}n-\frac{1}{2}\frac{\omega_{\lambda\nu}}{T}S^{\lambda\nu}\,,$
where $\rho=u_{\mu}u_{\nu}T^{\mu\nu}$ is the energy density, $n=u_{\mu}j^{\mu}$ is the charge density, and $S^{\lambda\nu}=u_{\mu}{\cal S}^{\mu\lambda\nu}$ is the spin density. Thus, defining the pressure $p$ as $T\phi\cdot u$, we have
$Ts=p+\rho-\mu n-\frac{1}{2}\omega_{\lambda\nu}S^{\lambda\nu}\,,$ (30)
which is the first thermodynamic relation in eq. (5). Following this definition and using the eq. (28) we have:
$p$ $=$ $T\phi^{\mu}u_{\mu}\,,$
$=$ $T\int^{T}_{0}\frac{{\rm d}T^{\prime}}{{T^{\prime}}^{2}}\Big{(}T^{\mu\nu}[T^{\prime}]u_{\nu}u_{\mu}-\mu u_{\mu}j^{\mu}[T^{\prime}]-$
-- $\frac{1}{2}\omega_{\lambda\nu}u_{\mu}\mathcal{S}^{\mu\lambda\nu}[T^{\prime}]\Big{)}$
$=$ $T\int^{T}_{0}\frac{{\rm d}T^{\prime}}{{T^{\prime}}^{2}}\left(\rho[T^{\prime}]-\mu n[T^{\prime}]-\frac{1}{2}\omega_{\lambda\nu}S^{\lambda\nu}[T^{\prime}]\right)\,.$
While the above equation, alongside with the (30), implies:
$\frac{\partial p}{\partial T}\Big{|}_{\mu,\omega}=s\,,$
the equations (7) and (8) cannot be derived from the equation (III). In fact, we will show in the following that for a system of massless free fermions and massive free fermions at global thermodynamic equilibrium with rotation and acceleration, the differential relation (8) does not hold.
III.1 Massless Dirac fermions
The exact mean values of the currents normalized so as to vanish in the Minkowski vacuum $|0_{M}\rangle$, at global thermodynamic equilibrium with the density operator (17) have been derived in ref. [34, 35] for the massless Dirac field. In order not to confuse them with the mean values normalized so as to vanish in the lowest lying state of the operator $\widehat{\Upsilon}$, that is, in the limit $T\to 0$, we will denote them by a tilde, hence:
$\tilde{\cal X}={\rm Tr}({\widehat{\rho}}_{\rm GE}\widehat{\cal X})-\langle 0_{M}|\widehat{\cal X}|0_{M}\rangle\,.$
What we need is, however:
${\cal X}$ $={\rm Tr}({\widehat{\rho}}_{\rm GE}\widehat{\cal X})-\langle 0|\widehat{\cal X}|0\rangle$
$={\rm Tr}({\widehat{\rho}}_{\rm GE}\widehat{\cal X})-\langle 0_{M}|\widehat{\cal X}|0_{M}\rangle+\langle 0_{M}|\widehat{\cal X}|0_{M}\rangle-\langle 0|\widehat{\cal X}|0\rangle$
$=\tilde{\cal X}+\langle 0_{M}|\widehat{\cal X}|0_{M}\rangle-\langle 0|\widehat{\cal X}|0\rangle\,.$
From the above equation, making use of the (22), we have:
$\lim_{T\to 0}\tilde{\cal X}=-\langle 0_{M}|\widehat{\cal X}|0_{M}\rangle+\langle 0|\widehat{\cal X}|0\rangle\,,$
which leads to the useful identity:
${\cal X}=\tilde{\cal X}-\lim_{T\to 0}\tilde{\cal X}\,,$ (32)
providing us with a straightforward method to calculate the mean values appearing in the entropy equation. For instance, the vector and axial current normalized so as to vanish in the Minkowski vacuum read [36, 37, 38, 34, 35]:
$\tilde{j}^{\mu}$ $=$ $\frac{\zeta\,T}{\pi^{2}}\left(\frac{T^{2}\left(\pi^{2}+\zeta^{2}\right)}{3}-\frac{a^{2}}{4}+\frac{\omega^{2}}{4}\right)u^{\mu}-\frac{\zeta\,l^{\mu}\,T^{3}}{6\pi^{2}}\,,$
$\tilde{j}^{\mu}_{A}$ $=$ $\frac{T^{2}}{\pi^{2}}\left(\frac{\pi^{2}}{6}+\frac{\zeta^{2}}{2}-\frac{\omega^{2}}{24\,T^{2}}-\frac{a^{2}}{8\,T^{2}}\right)\omega^{\mu}\,,$ (33)
where $a^{\mu}$ and $\omega^{\mu}$ have been defined in the eq. (19) and:
$l^{\mu}=\epsilon^{\mu\nu\rho\sigma}w_{\nu}\alpha_{\rho}u_{\sigma}\,.$
Applying equation (32), we get the following simple forms for the vector and axial currents to be used in the entropy current calculation:
$j^{\mu}$ $=$ $\frac{1}{3}\mu T^{2}u^{\mu}\,,$
$j^{\mu}_{A}$ $=$ $\frac{1}{6}T^{2}\omega^{\mu}\,.$ (34)
In refs. [34, 35] only the exact expression of the symmetric Belinfante stress-energy tensor was derived. However, since we need to tackle thermodynamic relations with a spin tensor, we have to select a pseudo-gauge where the spin tensor does not vanish. For the free Dirac field, the simplest pseudo-gauge with a non-trivial spin tensor is the canonical one, which is related to the Belinfante pseudo-gauge by the following equation:
$T^{\mu\nu}_{\rm Can}=T^{\mu\nu}_{\rm B}-\frac{1}{2}\partial_{\lambda}\mathcal{S}^{\lambda\mu\nu}_{\rm Can}\,.$ (35)
where the subscript ‘Can’ represents canonical currents and ‘B’ stands for Belinfante. The mean value of the canonical spin tensor is related to the axial current:
$\mathcal{S}^{\lambda\mu\nu}_{\rm Can}=-\frac{1}{2}\epsilon^{\lambda\mu\nu\rho}j_{A\rho}=-\frac{1}{12}\epsilon^{\lambda\mu\nu\rho}T^{2}\omega_{\rho}\,,$ (36)
where we used the axial current (III.1). Now, contracting Eq. (35) with $u_{\mu}u_{\nu}$ and using the Belinfante energy-momentum tensor derived in [35] (see Appendix A) we obtain:
$u_{\mu}u_{\nu}T^{\mu\nu}_{\rm Can}=u_{\mu}u_{\nu}T^{\mu\nu}_{\rm B}-\frac{1}{2}u_{\mu}u_{\nu}\partial_{\lambda}\mathcal{S}^{\lambda\mu\nu}_{\rm Can}$ (37)
$=u_{\mu}u_{\nu}T^{\mu\nu}_{\rm B}=\rho=\frac{7\pi^{2}T^{4}}{60}+\frac{\mu^{2}T^{2}}{2}-\frac{a^{2}T^{2}}{24}-\frac{\omega^{2}T^{2}}{8},$
where we took advantage of the antisymmetry of the indices in (36). By using the equations (III.1) and (36) we can obtain the contractions:
$n$ $=u_{\mu}j^{\mu}=\frac{1}{3}\mu T^{2}\,,$ (38)
$\omega_{\mu\nu}S^{\mu\nu}$ $=-\frac{T^{2}}{12}\epsilon^{\lambda\mu\nu\rho}\omega_{\rho}\omega_{\mu\nu}u_{\lambda}\,,$
$=-\frac{T^{2}}{12}\epsilon^{\lambda\mu\nu\rho}\epsilon_{\mu\nu\alpha\beta}\omega_{\rho}\omega^{\alpha}u^{\beta}u_{\lambda}=-\frac{T^{2}}{6}\omega^{2}\,.$ (39)
With the equations (37), (38) and (39) we are now in a position to determine the pressure by integrating the right-hand side of equation (III). Splitting the pressure of equation (III) into three terms:
$p=p_{1}-p_{2}-\frac{1}{2}p_{3}\,,$ (40)
we get:
$p_{1}$ $=$ $T\int^{T}_{0}\frac{{\rm d}T^{\prime}}{{T^{\prime}}^{2}}\rho[T^{\prime}]$
$=$ $\frac{7\pi^{2}}{180}T^{4}+\frac{1}{2}\mu^{2}T^{2}-\frac{{a}^{2}}{24}T^{2}-\frac{{\omega}^{2}}{8}T^{2}\,,$
$p_{2}$ $=$ $T\int^{T}_{0}\frac{{\rm d}T^{\prime}}{{T^{\prime}}^{2}}\mu n[T^{\prime}]=\frac{1}{3}\mu^{2}T^{2}\,,$
$p_{3}$ $=$ $T\int^{T}_{0}\frac{{\rm d}T^{\prime}}{{T^{\prime}}^{2}}\omega_{\mu\nu}S^{\mu\nu}[T^{\prime}]=-\frac{1}{6}\omega^{2}T^{2}\,.$
By using the equations (III.1), the equation (40) yields the pressure:
$p=\frac{7\pi^{2}}{180}T^{4}+\frac{1}{6}\mu^{2}T^{2}-\left({a}^{2}+{\omega}^{2}\right)\frac{T^{2}}{24}\,.$ (42)
While the entropy density can be obtained from the equation (30) by using the (37) and (38):
$s=\frac{7\pi^{2}}{45}T^{3}+\frac{1}{3}\mu^{2}T-\left({a}^{2}+{\omega}^{2}\right)\frac{T}{12}\,.$
Now, from (42)
$\frac{\partial p}{\partial T}\Big{|}_{\mu,\omega}=\frac{7\pi^{2}{T}^{3}}{45}+\frac{1}{3}\mu^{2}T-\left({a}^{2}+{\omega}^{2}\right)\frac{T}{12}\equiv s\,,$ (43)
as expected, and:
$\frac{\partial p}{\partial\mu}\Big{|}_{T,\omega}=\frac{1}{3}\mu T^{2}\equiv n\,,$ (44)
which is also in agreement with the differential relation (5). Yet:
$\frac{\partial p}{\partial\omega_{\lambda\nu}}\Big{|}_{T,\mu}$ $=$ $\frac{\partial p}{\partial{a}^{2}}\frac{\partial{a}^{2}}{\partial\omega_{\lambda\nu}}+\frac{\partial p}{\partial{\omega}^{2}}\frac{\partial{\omega}^{2}}{\partial\omega_{\lambda\nu}}\,,$
$=$ $-\frac{T^{2}}{12}\left(a_{\rho}\frac{\partial a^{\rho}}{\partial\omega_{\lambda\nu}}+\omega_{\rho}\frac{\partial\omega^{\rho}}{\partial\omega_{\lambda\nu}}\right)\,,$
$=$ $-\frac{T^{2}}{12}\left({a}^{\rho}u^{\sigma}\delta^{\lambda\nu}_{\rho\sigma}-\frac{1}{2}\epsilon^{\rho\alpha\beta\gamma}{\omega}_{\rho}u_{\gamma}\delta^{\lambda\nu}_{\alpha\beta}\right)\,,$
$=$ $\frac{T^{2}}{12}\left({a}^{\nu}u^{\lambda}-{a}^{\lambda}u^{\nu}+\epsilon^{\lambda\nu\rho\gamma}\omega_{\rho}u_{\gamma}\right)\,,$
$=$ $\frac{T^{2}}{12}\left({a}^{\nu}u^{\lambda}-{a}^{\lambda}u^{\nu}\right)+S^{\lambda\nu}\,,$
where we have used the equation (36) and the definition of the spin density $S^{\lambda\nu}=u_{\mu}{\cal S}^{\mu\lambda\nu}_{\rm Can}$. Hence, the derivative of the pressure with respect to the spin potential does not match the spin density and the relation (8) does not apply.
III.2 Massive Dirac fermions
A similar calculation can be carried out for massive fermions at second order of the expansion in thermal vorticity at global thermodynamic equilibrium. The general expression, up to that order, of the stress-energy tensor, the vector and axial currents was obtained in ref. [39]:
$T^{\mu\nu}_{\rm B}$ $=$ $\left(\rho-\frac{a^{2}U_{\alpha}}{T^{2}}-\frac{\omega^{2}U_{w}}{T^{2}}\right)u^{\mu}u^{\nu}+\frac{A}{T^{2}}a^{\mu}a^{\nu}$ (46)
-- $\left(p-\frac{a^{2}D_{\alpha}}{T^{2}}-\frac{\omega^{2}D_{w}}{T^{2}}\right)\Delta^{\mu\nu}+\frac{W}{T^{2}}\omega^{\mu}\omega^{\nu}$
$+$ $G\left(u^{\mu}l^{\nu}+u^{\nu}l^{\mu}\right)\,,$
$j^{\mu}$ $=$ $n\,u^{\mu}+\left(\frac{a^{2}N_{\alpha}}{T^{2}}+\frac{\omega^{2}N_{w}}{T^{2}}\right)u^{\mu}+G_{V}l^{\mu}\,,$ (47)
$j^{\mu}_{A}$ $=$ $\frac{1}{T}\omega^{\mu}W^{A}\,.$ (48)
The expression of the coefficients $\rho$, $p$, $U_{\alpha}$, $U_{w}$, $A$, $D_{\alpha}$, $D_{w}$, $W$, $G$, $n$, $N_{\alpha}$, $N_{w}$, $G_{V}$, and $W_{A}$ is reported in Appendix B in the equation (62). By using the equations (35) and (36) we find:
$u_{\mu}u_{\nu}T^{\mu\nu}_{\rm Can}$ $=$ $\rho-\frac{a^{2}U_{\alpha}}{T^{2}}-\frac{\omega^{2}U_{w}}{T^{2}}\,,$
$\mu u_{\mu}j^{\mu}$ $=$ $\mu\,n+\frac{\mu}{T^{2}}\left(a^{2}N_{\alpha}+\omega^{2}N_{w}\right)\,,$
$u_{\mu}\omega_{\lambda\nu}\mathcal{S}^{\mu\lambda\nu}_{\rm Can}$ $=$ $-\frac{1}{T}\omega^{2}W^{A}\,.$ (49)
One can work out the coefficients in eq. (62), in the ultra-relativistic and the non-relativistic limits.
The ultra-relativistic limit $m\rightarrow 0$ is just a control test, because it should coincide with our previous calculation. Indeed, it can be shown that in this limit the coefficients in eq. (62) reduce to:
$\rho$ $=$ $\frac{7\pi^{2}T^{4}}{60}+\frac{\mu^{2}T^{2}}{2}+\frac{\mu^{4}}{4\pi^{2}},\quad U_{\alpha}=\left(\frac{T^{4}}{24}+\frac{\mu^{2}T^{2}}{8\pi^{2}}\right),$
$U_{w}$ $=$ $\frac{1}{8}\left(T^{4}+\frac{3\mu^{2}T^{2}}{\pi^{2}}\right)\,,\quad W^{A}=\frac{T^{3}}{6}+\frac{T\mu^{2}}{2\pi^{2}}\,,$ (50)
$n$ $=$ $\frac{T^{2}\mu}{3}\left(1+\frac{\mu^{2}}{\pi^{2}T^{2}}\right)\,,\quad N_{\alpha}=N_{w}=\frac{T^{2}\mu}{4\pi^{2}}\,,$
thus, eqs. (49) become:
$u_{\nu}u_{\mu}T^{\mu\nu}_{\rm Can}$ $=$ $\frac{7\pi^{2}T^{4}}{60}+\frac{\mu^{2}T^{2}}{2}-\frac{a^{2}}{24}T^{2}-\frac{\omega^{2}}{8}T^{2},$
$\mu u_{\mu}j^{\mu}$ $=$ $\frac{T^{2}\mu^{2}}{3},$
$u_{\mu}\omega_{\lambda\nu}\mathcal{S}^{\mu\lambda\nu}_{\rm Can}$ $=$ $-\omega^{2}\frac{T^{2}}{6}.$ (51)
After subtracting the limits for $T\to 0$, we can obtain the pressure function of eq. (III), which fully matches eq. (42).
After having successfully checked the ultra-relativistic limit, we are now in a position to study the non-relativistic case $m/T\gg 1$. Setting $x=m/T$, we use the asymptotic expansion of the McDonald functions as [39]
$K_{\nu}(nx)\simeq{\rm e}^{-nx}\sqrt{\frac{\pi}{2x}}\left[1+\frac{4\nu^{2}-1}{8nx}+\dots\right].$ (52)
Thus, retaining only the dominant contribution in the non-relativistic limit, with $\mu<m$ or $\zeta<x$ (see definition (13)) the coefficients in eq. (62) are well approximated by:
$\rho$ $=$ $\frac{m^{4}}{\sqrt{2}(\pi x)^{3/2}}{\rm e}^{-(x-\zeta)}\,,$
$U_{\alpha}$ $=$ $\frac{x\,m^{4}}{24\sqrt{2}(\pi x)^{3/2}}{\rm e}^{-(x-\zeta)}\,,$
$U_{w}$ $=$ $\frac{m^{4}}{8\sqrt{2}(\pi x)^{3/2}}{\rm e}^{-(x-\zeta)}\,,$
$W^{A}$ $=$ $\frac{m^{3}}{2\sqrt{2}(\pi x)^{3/2}}{\rm e}^{-(x-\zeta)}\,,$
$n$ $=$ $\frac{m^{3}}{\sqrt{2}(\pi x)^{3/2}}{\rm e}^{-(x-\zeta)}\,,$
$N_{\alpha}$ $=$ $\frac{x\,m^{3}}{24\sqrt{2}(\pi x)^{3/2}}{\rm e}^{-(x-\zeta)}\,,$
$N_{w}$ $=$ $\frac{m^{3}}{8\sqrt{2}(\pi x)^{3/2}}{\rm e}^{-(x-\zeta)}\,,$ (53)
so that the eqs. (49) become:
$u_{\nu}u_{\mu}T^{\mu\nu}_{\rm Can}$ $=$ $\frac{m^{4}{\rm e}^{-(x-\zeta)}}{\sqrt{2}(\pi x)^{3/2}}\left(1-\frac{a^{2}x}{24\,T^{2}}-\frac{\omega^{2}}{8\,T^{2}}\right),$
$\mu u_{\mu}j^{\mu}$ $=$ $\frac{\mu\,m^{3}{\rm e}^{-(x-\zeta)}}{\sqrt{2}(\pi x)^{3/2}}\left(1+\frac{a^{2}x}{24\,T^{2}}+\frac{\omega^{2}}{8\,T^{2}}\right),$
$u_{\mu}\omega_{\lambda\nu}\mathcal{S}^{\mu\lambda\nu}_{\rm Can}$ $=$ $-\frac{m^{3}\,\omega^{2}}{2\sqrt{2}\,T(\pi x)^{3/2}}{\rm e}^{-(x-\zeta)}.$ (54)
Therefore, in the non-relativistic limit, the pressure $p$ defined in eq. (III) reads:
$p$ $=$ $\frac{m{\rm e}^{-x}\sqrt{x}}{96\sqrt{2}\pi^{3/2}(m-\mu)^{3}}\big{(}-ma^{2}(\mu+m)\big{(}{\rm e}^{\zeta}\big{(}4(\mu^{2}+m^{2})$ (55)
-- $8\mu m+6T(m-\mu)-3T^{2}\big{)}+3T^{3/2}\sqrt{\pi(m-\mu)}{\rm e}^{x}\big{)}$
-- $6T(m-\mu)\big{(}{\rm e}^{\zeta}\big{(}\omega^{2}(\mu+m)(2(m-\mu)-T)$
$+$ $4(\mu-m)\omega^{2}T-64T(m-\mu)^{3}\big{)}$
$+$ $\sqrt{\pi T(m-\mu)}{\rm e}^{x}\big{(}\omega^{2}(\mu+m)+32(m-\mu)^{3}\big{)}\big{)}\big{)}\,.$
We can now evaluate its derivative with respect to $\omega$:
$\frac{\partial p}{\partial\omega_{\lambda\nu}}\Big{|}_{T,\mu}$ $=$ $\frac{\partial p}{\partial{a}^{2}}\frac{\partial{a}^{2}}{\partial\omega_{\lambda\nu}}+\frac{\partial p}{\partial{\omega}^{2}}\frac{\partial{\omega}^{2}}{\partial\omega_{\lambda\nu}}$
$=$ $2\frac{\partial p}{\partial{a}^{2}}\,a\frac{\partial a}{\partial\omega_{\lambda\nu}}+2\frac{\partial p}{\partial{\omega}^{2}}\,\omega\frac{\partial\omega}{\partial\omega_{\lambda\nu}}$
$=$ $2\frac{\partial p}{\partial{a}^{2}}{a}^{\rho}u^{\sigma}\delta^{\lambda\nu}_{\rho\sigma}+2\frac{\partial p}{\partial{\omega}^{2}}{\omega}_{\rho}\left(-\frac{1}{2}\epsilon^{\rho\alpha\beta\gamma}u_{\gamma}\right)\delta^{\lambda\nu}_{\alpha\beta}$
$=$ $2\frac{\partial p}{\partial{a}^{2}}\left({a}^{\lambda}u^{\nu}-{a}^{\nu}u^{\lambda}\right)-2\frac{\partial p}{\partial{\omega}^{2}}\epsilon^{\rho\lambda\nu\gamma}{\omega}_{\rho}u_{\gamma}\,,$
where:
$\frac{\partial p}{\partial{a}^{2}}$ $=$ $-\frac{m^{5/2}(m+\mu)}{96\pi^{3/2}\sqrt{2T}(m-\mu)^{3}}\Big{(}{\rm e}^{(\zeta-x)}\big{(}4\left(\mu^{2}+m^{2}\right)$
-- $8\mu m+6T(m-\mu)-3T^{2}\big{)}+3\sqrt{\pi}T^{3/2}\sqrt{m-\mu}\Big{)},$
$\frac{\partial p}{\partial{\omega}^{2}}$ $=$ $-\frac{m\sqrt{mT}}{16\sqrt{2}\pi^{3/2}(m-\mu)^{2}}\big{(}{\rm e}^{(\zeta-x)}\big{(}2\left(m^{2}-\mu^{2}\right)$ (57)
-- $5mT+3\mu T\big{)}+(\mu+m)\sqrt{\pi T(m-\mu)}\big{)}.$
The canonical spin density may be obtained like in the equation (36) by using the (48) and the approximated expression in (III.2) for $W_{A}$:
$S^{\mu\nu}$ $=$ $u_{\lambda}\mathcal{S}^{\lambda\mu\nu}_{\rm Can}=-\frac{1}{2}\epsilon^{\lambda\mu\nu\rho}u_{\lambda}j_{A\rho}=-\frac{1}{2T}\epsilon^{\lambda\mu\nu\rho}W_{A}\omega_{\rho}u_{\lambda}$ (58)
$\simeq$ $-\frac{1}{2T}\epsilon^{\lambda\mu\nu\rho}\omega_{\rho}u_{\lambda}\frac{m^{3}}{2\sqrt{2}(\pi x)^{3/2}}{\rm e}^{-(x-\zeta)}\,.$
Comparing the (58) with the (III.2)-(III.2) we can readily realize that:
$\frac{\partial p}{\partial\omega_{\lambda\nu}}\Big{|}_{T,\mu}\neq S^{\lambda\nu}\,.$ (59)
Again, like in the massless case, in the non-relativistic limit the derivative of the pressure has an additional component which is proportional to $(a^{\lambda}u^{\nu}-a^{\nu}u^{\lambda})$ of the same order of magnitude as the one proportional to the vorticity. Hence, the same conclusion holds.
IV Conclusions
In conclusion, we have shown, by providing two specific counter-examples, namely a system of massless free fermions and massive free fermions with rotation and acceleration at global thermodynamic equilibrium, that the differential thermodynamic relations (6) often used to obtain the constitutive equations in relativistic spin hydrodynamics do not hold. Notably, the derivative of the pressure function with respect to spin potential (which reduces to vorticity at global equilibrium) features additional contributions of the same order of magnitude as the spin density. The relations (5) and (6) therefore can hardly be used to make a precise derivation of the constitutive relations in relativistic spin hydrodynamics; their use may easily lead to missing some dissipative terms for the spin tensor. Indeed, the quantum-statistical method proposed in refs. [26, 32] to determine the entropy production rate and constitutive equations in relativistic spin hydrodynamics led to the identification of several new dissipative coefficients for the spin tensor.
Acknowledgements.
R.S. acknowledges the kind hospitality and support of INFN Firenze, Department of Physics and Astronomy, University of Florence, and ECT* Trento, where part of this work has been carried out. R.S. is also supported partly by a postdoctoral fellowship of West University of Timișoara, Romania. We thank the Galileo Galilei Institute for Theoretical Physics for the hospitality and the INFN for partial support during the completion of this work.
Appendix A Energy-momentum tensor for free massless Dirac field
For the case of global equilibrium with rotation and acceleration, we have the general Belinfante form of energy-momentum tensor for massless fermions as [34, 35]
$T^{\mu\nu}_{\rm B}$ $=$ $\rho u^{\mu}u^{\nu}-p\Delta^{\mu\nu}+Ww^{\mu}w^{\nu}+A\alpha^{\mu}\alpha^{\nu}+G^{l}l^{\mu}l^{\nu}$ (60)
$+$ $G(l^{\mu}u^{\nu}+l^{\nu}u^{\mu})+\mathbb{A}(\alpha^{\mu}u^{\nu}+\alpha^{\nu}u^{\mu})$
$+$ $G^{\alpha}(l^{\mu}\alpha^{\nu}+l^{\nu}\alpha^{\mu})+\mathbb{W}(w^{\mu}u^{\nu}+w^{\nu}u^{\mu})$
$+$ $A^{w}(\alpha^{\mu}w^{\nu}+\alpha^{\nu}w^{\mu})+G^{w}(l^{\mu}w^{\nu}+l^{\nu}w^{\mu})\,,$
where
$\rho$ $=$ $T^{4}\frac{7\pi^{4}+30\pi^{2}\zeta^{2}+15\zeta^{4}}{60\pi^{2}}-\frac{a^{2}T^{2}(\pi^{2}+3\zeta^{2})}{24\pi^{2}}-\frac{17a^{4}}{960\pi^{2}}$
-- $\frac{\omega^{2}T^{2}(\pi^{2}+3\zeta^{2})}{8\pi^{2}}+\frac{23\,a^{2}\,\omega^{2}}{1440\pi^{2}}+\frac{\omega^{4}}{64\pi^{2}}+\frac{11(a\cdot\omega)^{2}}{720\pi^{2}},$
$G$ $=$ $T^{4}\frac{\pi^{2}+3\zeta^{2}}{18\pi^{2}}-\frac{31\,a^{2}T^{2}}{360\pi^{2}}-\frac{13\,\omega^{2}T^{2}}{120\pi^{2}},$ (61)
$W$ $=$ $-\frac{a^{2}T^{2}}{27\pi^{2}},\quad A=-\frac{\omega^{2}T^{2}}{27\pi^{2}},\quad A^{w}=\frac{a\cdot\omega\,T^{2}}{27\pi^{2}},$
$p$ $=$ $\frac{\rho}{3},\quad G^{l}=-\frac{2\,T^{4}}{27\pi^{2}},\quad\mathbb{A}=\mathbb{W}=G^{\alpha}=G^{w}=0.$
Appendix B Coefficients of the energy-momentum tensor for free massive Dirac field
The coefficients $\rho$, $U_{\alpha}$, $U_{w}$, $n$, $N_{\alpha}$, $N_{w}$, $A$, $W$, $p$, $D_{\alpha}$, $D_{w}$, $G$, $G_{V}$, and $W^{A}$ of the stress-energy tensor, the vector and axial currents for the massive Dirac fields are given by the general expressions [39]:
$\rho$ $=$ $-\frac{2m^{4}}{\pi^{2}}\Sigma\left[\frac{K_{3}(nx)}{nx}-\frac{K_{2}(nx)}{(nx)^{2}}\right]\cosh{(n\,\zeta)},$
$p$ $=$ $-\frac{2m^{4}}{\pi^{2}}\,\Sigma\left[\frac{K_{2}(nx)}{(nx)^{2}}\right]\cosh(n\,\zeta),$
$U_{\alpha}$ $=$ $-\frac{m^{4}}{12\pi^{2}}\Sigma\left[(3x^{-2}+n^{2})K_{2}(nx)\right]\cosh{(n\,\zeta)},$
$D_{\alpha}$ $=$ $\frac{m^{4}}{12\pi^{2}}\Sigma\left[\frac{3K_{2}(nx)}{x^{2}}-n\frac{K_{3}(nx)}{x}\right]\cosh(n\,\zeta),$
$A$ $=$ $W=0,$
$D_{w}$ $=$ $-\frac{m^{4}}{4\pi^{2}}\Sigma\left[\frac{K_{2}(nx)}{x^{2}}\right]\cosh(n\,\zeta),$
$U_{w}$ $=$ $-\frac{m^{4}}{4\pi^{2}}\Sigma\left[-\frac{K_{2}(nx)}{x^{2}}+n\frac{K_{3}(nx)}{x}\right]\cosh{(n\,\zeta)},$
$W^{A}$ $=$ $-\frac{m^{3}}{\pi^{2}}\Sigma\left[\frac{K_{0}(nx)}{x}+\frac{2K_{1}(nx)}{nx^{2}}\right]\cosh{(n\,\zeta)},$
$n$ $=$ $-\frac{2m^{3}}{\pi^{2}}\Sigma\left[\frac{K_{2}(nx)}{nx}\right]\sinh{(n\,\zeta)},$
$N_{\alpha}$ $=$ $-\frac{m^{3}}{12\pi^{2}}\Sigma\left[n^{2}K_{1}(nx)+\frac{3nK_{2}(nx)}{x}\right]\sinh{(n\,\zeta)},$
$N_{w}$ $=$ $-\frac{m^{3}}{4\pi^{2}}\Sigma\left[\frac{2nK_{0}(nx)}{x^{2}}+\frac{K_{1}(nx)}{x}\right]\sinh{(n\,\zeta)},$
$G$ $=$ $-\frac{m^{4}}{12\pi^{4}}\Sigma\left[n\frac{K_{3}(nx)}{x}\right]\cosh(n\,\zeta),$
$G_{V}$ $=$ $\frac{m^{3}}{6\pi^{2}}\Sigma\left[\frac{2nK_{0}(nx)}{x^{2}}+\frac{K_{1}(nx)}{x}\right]\sinh{(n\,\zeta)},$ (62)
where $\Sigma\equiv\Sigma_{n=1}^{\infty}\left(-1\right)^{n+1}$, $\zeta=\mu/T$, and $x=m/T$.